OpenStudy (anonymous):

And i need help with these problem. An alloy of tin is 15% tin and weighs 20 pounds. A second alloy is 10% tin. How much (to the nearest tenth lb.) of the second alloy must be added to the first alloy to get a 12% mixture. ______ pounds

OpenStudy (anonymous):

Problem 1: A jeweler is using an alloy, a mixture of metals. He has 10 pounds of the alloy. 2% is the rare metal robium. He has another source which is 5% robium. He wishes to combine these mixtures to form a mixture that is 3% robium. Solution: 1. Identify the variables. (0.02)(10) The quantity of robium in material 1 (5%)(x) The quantity of robium in material 2 (0.03)(x+10) The quantity of the desired mixture 2. Write an equation. 5% x + (0.02)(10) = (0.03)(x+10) The equation required. 3. Perform the Solution. 0.05x + 0.2 = .03x + 0.3 = 0.02x = 0.1 4. Write the answer. x = 5 The desired amount.

OpenStudy (andriod09):

just do what i did. use the percent bar, then the butterfly method. want me to start for you? here: